Consider any two random variables X, Y of any distirbution and not necesarily independent. Given that P(X=a)=p1, P(max(X, Y) =a) =p2, and P(min(X, Y) =a) =p3.

Find P(Y=a) in terms of p1, p2 and p3.

Hint: Use the first Bayes theorem. Choose a suitable event-complement pair to partition the probability space by comparing X and Y .
Answer #1


Complete solution for the problem can be found in the images below with all the necessary steps.
Do give a thumbs up if you liked the solution. Thank You :)こ P (Y= a) P(Y=a/xsy). PCX5Y) +P [ Y=0 / x70Y) -P(x>Y) P(Y Man(X,Y)= ) P( mon(x,y)= a) = P(x=a ,Y<a) + P(Y=a, x < a) +1(x=a,(1L we Adding 0 s get, Р{~~[9, 1)- а) + (-m(x,)) (а, ч<b) + ( Yen, xce ) +P (у-а,154) +Р (Y-а , X}a +24 Р Х = 4, Y = 4 а). х7 P(x=a, Y>a) + P(x=a Y=a) + P(x=a, y <a pl L (un Noe Z IV . > PLY= a) P (Y= a, xya) + P (Y=a. X<a a) + P ( 4 = a, X=a) from7. PC PC 7- а Ya ) = P2 + F3 21.